OpenStudy (anonymous):

Which absolute value function is equivalent to the piecewise function shown ? A) f(x) = |x| - 2 B) f(x) = |x - 2| C) f(x) = |x| + 2 D) f(x) = |x + 2|

OpenStudy (zarkon):

what piecewise function shown ?

OpenStudy (anonymous):

huh

OpenStudy (zarkon):

\[\text{Which absolute value function is equivalent to the }\underline{\text{piecewise function shown}} ?\] I see no piecewise function (excluding the absolute value functions)

OpenStudy (anonymous):

http://usatestprep.com/modules/gallery/files/52/5212/5212.JPG go to this website ..

OpenStudy (zarkon):

D

OpenStudy (anonymous):

Which shows f(x) = |3x + 3| in piecewise form? A) B) C) D)

OpenStudy (anonymous):

http://usatestprep.com/modules/gallery/files/25/2543/2543.jpg here is the website thinggyy

OpenStudy (zarkon):

\[f(x)=\left\{\begin{matrix}3x + 3& \text{if} & x\ge -1\\ -(3x + 3)& \text{if} &x<-1\end{matrix}\right.\] \[=\left\{\begin{matrix}3x + 3& \text{if} & x\ge -1\\ -3x - 3& \text{if} &x<-1\end{matrix}\right.\]

OpenStudy (anonymous):

website thinggy! i like it

OpenStudy (anonymous):

hahah yeeeepppp